# Panoramic Distortion-Aware Tokenization for Person Detection and Localization Using Transformers in Overhead Fisheye Images

### v1

@@ 4 Experiments
To demonstrate the validity and effectiveness of our approach, we conducted extensive experiments using a large-scale dataset, LOAF [63], composed of overhead fisheye images in Table 2.
Table 2: Properties of the LOAF dataset
Dataset Resolution Number of images
Train Validation Test All
LOAF [63] 1024 $\times$ 1024 29,569 4600 8773 42,942
@@
Method val test
mAP AP50 AP75 APn APm APf APseen APunseen mAP AP50 AP75 APn APm APf APseen APunseen
Seidel et al. [40] VISAPP’19 21.8 59.8 07.6 32.8 28.5 02.3 23.2 19.5 20.2 58.2 07.1 32.2 28.3 03.9 22.4 18.9
Li et al. [21] AVSS’19 28.5 63.3 20.1 46.8 24.2 01.3 33.8 29.3 27.2 65.2 21.3 47.2 24.8 01.3 31.8 27.6
  Tamura et al. [46] WACV’19 34.8 72.1 27.7 51.7 38.8 08.7 38.8 32.9 34.2 72.8 28.7 53.7 37.3 06.5 39.5 33.2
RAPiD [9] CVPRW’20 40.3 77.9 34.8 55.3 41.9 09.2 44.7 37.6 39.2 77.9 35.4 54.8 40.1 07.9 44.2 37.3
OARPD [33] MTA’24 45.5 78.8 46.3 - - - - - 45.8 79.4 45.8 - - - - -
Yang et al. [63] ICCV’23 47.4 82.6 48.4 64.1 54.3 14.1 50.8 45.7 46.2 81.1 47.3 66.1 53.5 12.6 49.3 44.9
Ours 52.7 84.9 57.1 74.8 51.6 15.3 56.8 50.9 55.2 87.6 61.2 71.1 51.8 09.8 56.8 54.2
@@ 4.1 Parameter settings
For our network backbone, we used SG-Former-S [36] (where the backbone size is the smallest among the S, M, and B types), which was pretrained on ImageNet [37]. We trained our network using a mini-batch size of 8 and the AdamW optimizer [29]. The initial learning rate was set at $2\times 10^{-4}$ and was multiplied by 0.1 at the 40th epoch. We used the hyper-parameters of $K=5$, $a=3$, and $\alpha=2$. The remaining hyper-parameters were applied in [63]. Unless otherwise specified, panoramic images with 3072 $\times$ 768 pixels were used.
@@ 4.2 Experimental results
We used the results of person detection and localization results reported by Yang et al. [63] in the methods of Seidel et al. [40], Li et al. [21], Tamura et al. [46], RAPiD [9], and Yang et al. [63]. We used the OARPD results reported in [33]. To acquire qualitative results, we used Yang et al.’s [63] official codes using PyTorch [31].
4.2.1 Person detection
@@
To demonstrate the validity and the effectiveness of the proposed method, we used the mean average precision (mAP) following a COCO evaluation [23]. Our method achieved the highest mAP among the methods listed in Table 3 on both the validation and test sets. The mAP in our method is substantially higher than that obtained using Yang et al.’s method [63] by values of 5.3 and 9.0 on the LOAF validation and test sets, respectively. Our method also achieved an APseen of 56.8 and an APunseen of 54.2 on the LOAF test set, thus outperforming the other methods. The substantial margin between our method and Yang et al.’s method [63] on the LOAF unseen test set was 9.3, which demonstrates scene robustness. In addition, our method outperformed the other methods on APn and our method demonstrated a performance comparable to that of Yang et al.’s method [63] on APm and APf.
@@ 4.2.2 Diagnostic study
Table 4: Ablation study on the LOAF test set
Method test
mAP AP50 AP75 APseen APunseen
Baseline (DAB-DETR) 40.2 74.0 39.6 42.5 38.7
$+$ Panoramic conversion 50.6 85.2 54.1 52.1 49.8
$+$ PDAT 55.2 87.6 61.2 56.8 54.2
Table 5: Panoramic-image resolution analysis on the LOAF test set
Panorama width test
mAP AP50 AP75 APseen APunseen
2048 pixel 46.9 81.5 47.9 47.9 46.2
2560 pixel 49.7 84.5 52.3 51.1 49.0
3072 pixel 55.2 87.6 61.2 56.8 54.2
We conducted a diagnostic study on the LOAF test set. Table 4 shows the ablation study results. The baseline indicates DAB-DETR [26] with the SG-Former-S [36] for the backbones using fisheye input images. The second row demonstrates the effectiveness of the panoramic conversion. The panorama images were suitable for DAB-DETR because they removed person rotations. After adopting the panoramic conversion, the performance increased by an mAP of 10.4. We validated the effectiveness of the PDAT process. The third row indicates that our method using PDAT provided notably improved person detection results when compared with our method without the PDAT by an mAP of 4.6. Therefore, the PDAT alleviated the problems caused by the variations in person size dramatically. Similar to mAP, both APseen and APunseen were improved notably. In addition, Table 5 shows the dependency of the panoramic image width. The 3072-pixel panoramic image width is suitable for the 1024$\times$1024-pixel fisheye images because the circumference of the fisheye image is $\pi\times 1024=3217\approx 3072$. Therefore, the 3072-pixel image width without shrinkage led to the best performance among 2048, 2560, and 3072 pixels.
Table 6: Person localization results using position errors (PE) $\downarrow$ in meters on the LOAF validation (val) and test sets
Method val test
   mPE PEn PEm PEf PEseen PEunseen    mPE PEn PEm PEf PEseen PEunseen
Seidel et al. [40] VISAPP’19 1.298 0.561 1.332 3.109 1.206 1.382 1.321 0.706 1.309 3.482 1.306 1.386
Li et al. [21] AVSS’19 0.898 0.502 0.871 2.650 0.832 0.962 0.913 0.543 0.884 2.780 0.904 0.998
  Tamura et al. [46] WACV’19 0.755 0.429 0.736 1.862 0.709 0.821 0.778 0.471 0.826 2.160 0.724 0.836
RAPiD [9] CVPRW’20 0.674 0.426 0.623 1.403 0.625 0.757 0.682 0.461 0.664 1.445 0.638 0.776
OARPD [33] MTA’24 - - - - - - - - - - - -
Yang et al. [63] ICCV’23 0.368 0.144 0.363 0.775 0.318 0.402 0.374 0.148 0.375 0.819 0.326 0.398
Ours 0.279 0.096 0.201 0.576 0.2695 0.306 0.305 0.117 0.253 0.598 0.345 0.276
Figure 5: Qualitative results on the LOAF test set. Green and red rectangles indicate ground-truth and estimated bounding boxes, respectively. (a) Our detection result in a panoramic image. (b) From left to right: detection results of Yang et al. [63] and our method. For visualization, rotated rectangles are used for our method.
@@ 4.2.3 Person localization
To validate the person localization accuracy, we compared our method with the conventional methods. We evaluated the individual person positions from the bounding boxes using Yang et al.’s approach [63], as described in Section 2. We defined the position error (PE) in meters as the horizontal distance between the ground-truth person position and the person position estimated from the bounding boxes. Our method achieved the lowest mean PE (mPE), as indicated in Table 6. The mPE for our method is smaller than that of Yang et al.’s method [63] on the LOAF test set by 0.069. Similar to the person detection performance, our method realized a noteworthy 0.122 gain in terms of the PEunseen when compared with Yang et al.’s method [63]. Our method outperformed the other methods on PEn, PEm, and PEf because of accurate bounding box detection.
4.2.4 Qualitative evaluation
To evaluate the detection accuracy, we compared the bounding boxes. Figure 5(a) shows our detection results in a panoramic image on the LOAF test set. Although people at the top of the panoramic image are small, our networks can accurately detect bounding boxes. We found that some tiny persons without ground-truth annotations were also detected, as shown in the zoom-in area of Figure 5(a). Our network has the potential to detect more persons. Overall, our results are the closest to the ground-truth bounding boxes. Furthermore, the transformer-based methods proposed by Yang et al. [63] failed to detect some people at near image circles, as shown in Figure 5(b).
@@ 5 Conclusion
Limitations. We assume that the fisheye images are a stereographic projections. However, this assumption may cause a mismatch with the actual distortion. To address this problem, we will extend our method using a deep single image camera calibration technique [4, 53] that estimates the lens distortion. Another promising direction for future work is to use videos for the input. In this paper, we have focused on person detection and localization from an image.

### v2

@@ 4 Experiments
To demonstrate the validity and effectiveness of our approach, we conducted extensive experiments using three large-scale datasets [10, 48, 64], which are composed of overhead fisheye images. We used 1024$\times$1024 input resolution by following the settings given in [10, 22, 41, 47, 64]. For evaluation, we converted the trapezoid bounding boxes into rectangles because no other methods listed in Table 1 produced trapezoid outputs.
@@
Method mAP AP50 AP75 APn APm APf APseen APunseen
val
Seidel et al. [41] VISAPP’19 21.8 59.8 07.6 32.8 28.5 02.3 23.2 19.5
Li et al. [22] AVSS’19 28.5 63.3 20.1 46.8 24.2 01.3 33.8 29.3
Tamura et al. [47] WACV’19 34.8 72.1 27.7 51.7 38.8 08.7 38.8 32.9
RAPiD [10]   CVPRW’20 40.3 77.9 34.8 55.3 41.9 09.2 44.7 37.6
OARPD [34] MTA’24 45.5 78.8 46.3 - - - - -
Yang et al. [64] ICCV’23 47.4 82.6 48.4 64.1 54.3 14.1 50.8 45.7
Ours 52.4 84.6 57.6 66.3 57.9 42.9 56.3 50.7
test
Seidel et al. [41] VISAPP’19 20.2 58.2 07.1 32.2 28.3 03.9 22.4 18.9
Li et al. [22] AVSS’19 27.2 65.2 21.3 47.2 24.8 01.3 31.8 27.6
Tamura et al. [47] WACV’19 34.2 72.8 28.7 53.7 37.3 06.5 39.5 33.2
RAPiD [10]   CVPRW’20 39.2 77.9 35.4 54.8 40.1 07.9 44.2 37.3
OARPD [34] MTA’24 45.8 79.4 45.8 - - - - -
Yang et al. [64] ICCV’23 46.2 81.1 47.3 66.1 53.5 12.6 49.3 44.9
Ours 54.5 87.2 61.1 62.8 60.5 44.6 56.2 53.6
@@ 4.1 Parameter settings
For our network backbone, we used SG-Former-S [37], which has the smallest backbone size among the S, M, and B types, and was pretrained on ImageNet [38]. We trained our network using a combination of a mini-batch size of 8 for 120 epochs and the AdamW optimizer [30]. The initial learning rate was set at $2\times 10^{-4}$ and was multiplied by 0.1 at the 40th epoch. We used the hyperparameters $K=5$ and $\alpha=2$. Because the majority of bounding boxes were near the image circles of fisheye images, we used the 3072-pixel width of panoramic images, approximately corresponding to the circumference of fisheye images with 1024 $\times$ 1024 pixels: $\pi\times 1024=3217\approx 3072$ pixels. Unless otherwise specified, we used panoramic images with 3072 $\times$ 768 pixels.
@@ 4.2 Experimental results
We used the person detection and localization results reported by Yang et al. [64] in the methods of Seidel et al. [41], Li et al. [22], Tamura et al. [47], RAPiD [10], and Yang et al. [64]. We also used the OARPD results reported in [34]. Unless otherwise specified, we refer to these results throughout. Because of the lack of previous performance reports, we implemented methods for comparison of Tamura et al. [47] and OARPD [34], based on the corresponding papers using PyTorch [32]. In addition, the official codes from Yang et al. [64] and RAPiD [10] [^1] Yang et al.: https://github.com/BUPT-PRIV/LOAF,  RAPiD: https://github.com/duanzhiihao/RAPiD were used for evaluation.
4.2.1 Person detection on LOAF
Table 3: Person detection results on the CEPDOF and WEPDTOF datasets
Method CEPDOF WEPDTOF
mAP AP50 AP75 Precision Recall F1-score mAP AP50 AP75 Precision Recall F1-score
Seidel et al. [41] 20.9 50.6 10.2 80.6 39.5 53.0 16.1 39.4 09.0 70.9 38.6 50.0
Li et al. [22] 34.2 75.7 28.6 86.3 65.4 74.4 25.2 69.9 30.2 81.4 64.5 72.0
Tamura et al. [47] 1 28.8 61.6 23.0 73.1 65.8 69.3 12.8 27.2 10.9 53.0 34.3 41.6
RAPiD [10] 1 38.3 78.6 30.8 91.8 80.3 85.7 25.0 54.0 17.4 83.7 57.6 68.2
OARPD [34] 1 11.4 28.9 06.6 80.5 31.3 45.1 04.1 10.0 02.2 28.8 19.0 22.9
Yang et al. [64] 1 39.7 78.3 37.1 82.9 72.6 77.4 20.5 45.3 16.3 66.8 43.0 52.3
Ours 42.8 85.6 38.7 91.8 81.8 86.5 37.5 74.9 34.8 81.7 69.0 74.8
1 We trained networks for evaluation.
Table 4: Ablation study on the LOAF test set
Method mAP AP50 AP75 APn APm APf APseen APunseen
Baseline (DAB-DETR) 40.2 74.0 39.6 48.2 49.5 32.7 42.5 38.7
$+$ Panoramic conversion 49.8 84.7 52.7 59.3 55.3 39.7 51.0 48.9
$+$ PDAT 54.5 87.2 61.1 62.8 60.5 44.6 56.2 53.6
Table 5: Panoramic-image resolution analysis on the LOAF test set
 Panoramic image width mAP AP50 AP75 APn APm APf APseen APunseen
2048 pixels 47.0 81.3 49.2 58.3 52.0 35.5 48.5 46.0
2560 pixels 51.2 85.1 55.6 61.9 56.4 39.3 53.8 49.7
3072 pixels 54.5 87.2 61.1 62.8 60.5 44.6 56.2 53.6
Metrics and protocols. To demonstrate the validity and the effectiveness of the proposed method on LOAF, we used the mean average precision (mAP) following a COCO evaluation [24]. We used the subscripts $n$, $m$, and $f$ to denote the near (0–10 m), mid-range (10–20 m), and distant (over 20 m) person positions, respectively, e.g., APn. LOAF also provides seen and unseen scenes for test splits using the subscripts $seen$ and $unseen$, respectively. We used the official splits of the training, validation, and test sets.
@@ 4.2.1 Person detection on LOAF
Results. Our method achieved the highest mAP among the methods listed in Table 2 when used on both the validation and test sets. The mAP obtained using our method was substantially higher than that obtained using Yang et al.’s method [64] by values of 5.0 and 8.3 on the LOAF validation and test sets, respectively. In addition, the APf value obtained using our method on the test set was notably larger than that of Yang et al.’s method [64] by 32.0, where 67% of the bounding boxes were labeled as $f$: over 20 m. Our method also achieved an APseen of 56.2 and an APunseen of 53.6 on the test set, thus outperforming the other methods. The substantial margin between our method and Yang et al.’s method [64] on the unseen test set of 8.7 demonstrates the scene robustness.
4.2.2 Person detection on CEPDOF
Metrics and protocols. CEPDOF [10] provides bounding box annotations without person locations in the world coordinates. Therefore, we evaluated the detection performance using the AP. Similar to prior works [10, 48, 64], we also used precision, recall, and the F1-score. Following the procedure of [64], we trained the method of Tamura et al. [47], along with RAPiD [10], OARPD [34], Yang et al. [64], and our method, using HABBOF [22] and WEPDTOF [48] after pretraining these methods on COCO [24], although MW-R [10] was not used for the training because it is not publicly available.
Results. Our method achieved the highest mAP, AP50, and AP75 values on CEPDOF among the methods listed in Table 3. The mAP, AP50, and AP75 values achieved using our method are higher than those obtained using Yang et al.’s method [64] by values of 3.1, 7.3, and 1.6, respectively. Additionally, the F1-score of our method was greater than that of Yang et al.’s method [64] by 9.1.
4.2.3 Person detection on WEPDTOF
Metrics and protocols. We used the evaluation metrics of CEPDOF, as described in Section 4.2.2. We trained the methods of Tamura et al. [47], RAPiD [10], OARPD [34], Yang et al. [64], and our method, using HABBOF and CEPDOF after they were pretrained on COCO following [64], although MW-R was again not used for the training because it is not publicly available.
Results. Similar to the CEPDOF results, our method achieved the highest mAP, AP50, and AP75 results on WEPDTOF among the methods listed in Table 3. The mAP values achieved on CEPDOF using both the conventional methods and our method were larger than those obtained on WEPDTOF because WEPDTOF has more scenes; i.e., WEPDTOF consists of 16 videos, and CEPDOF consists of 8 videos. Comparison between the CEPDOF and WEPDTOF results shows that the methods of Tamura et al. [47], RAPiD [10], OARPD [34], and Yang et al. [64] apparently require diverse images because the network is trained on fisheye images that include rotated persons.
@@ 4.2.4 Diagnostic study
We conducted a diagnostic study on the LOAF test set. Table 4 shows the ablation study results. The baseline represents DAB-DETR [27] with the SG-Former-S [37] for the backbone when using fisheye input images. The second row demonstrates the effectiveness of the panoramic conversion. The panoramic images were suitable for DAB-DETR because they removed the person rotations. After the panoramic conversion was adopted, the performance increased by an mAP of 9.6. We also validated the effectiveness of the PDAT process. The third row indicates that our method using PDAT provided substantially improved person detection results when compared with our method without PDAT by an mAP of 4.7. Therefore, the PDAT alleviated the problems caused by the variations in person size dramatically. Similar to the mAP, the other metrics were improved notably. In addition, Table 5 shows the dependence of the panoramic image width. The 3072-pixel panoramic image width is suitable for the 1024$\times$1024-pixel fisheye images because the circumference of the fisheye image is $\pi\times 1024=3217\approx 3072$. Therefore, the 3072-pixel image width without shrinkage led to the best performance among the widths of 2048, 2560, and 3072 pixels.
Table 6: Person localization results using position errors (PE) $\downarrow$ in meters on the LOAF val and test sets
Method    mPE PEn PEm PEf PEseen PEunseen
val
Seidel et al. [41] VISAPP’19 1.298 0.561 1.332 3.109 1.206 1.382
Li et al. [22] AVSS’19 0.898 0.502 0.871 2.650 0.832 0.962
  Tamura et al. [47] WACV’19 0.755 0.429 0.736 1.862 0.709 0.821
RAPiD [10] CVPRW’20 0.674 0.426 0.623 1.403 0.625 0.757
OARPD [34] MTA’24 - - - - - -
Yang et al. [64] ICCV’23 0.368 0.144 0.363 0.775 0.318 0.402
Ours 0.286 0.096 0.182 0.635 0.256 0.306
test
Seidel et al. [41] VISAPP’19 1.321 0.706 1.309 3.482 1.306 1.386
Li et al. [22] AVSS’19 0.913 0.543 0.884 2.780 0.904 0.998
  Tamura et al. [47] WACV’19 0.778 0.471 0.826 2.160 0.724 0.836
RAPiD [10] CVPRW’20 0.682 0.461 0.664 1.445 0.638 0.776
OARPD [34] MTA’24 - - - - - -
Yang et al. [64] ICCV’23 0.374 0.148 0.375 0.819 0.326 0.398
Ours 0.277 0.115 0.261 0.528 0.282 0.274
@@ 4.2.5 Person localization
Metrics and protocols. To validate the person localization accuracy, we compared our method with the conventional methods. We evaluated the individual person positions from the bounding boxes using Yang et al.’s approach [64], as described in Section 2. We used the position error (PE) in meters as the horizontal distance between the ground-truth person position and the person position estimated from the bounding boxes. We also used the same subscripts used in Section 4.2.1, e.g., PEn. These estimated bounding boxes were the same as the person detection results on LOAF from Section 4.2.1.
Results. Our method achieved the lowest mean PE (mPE), as indicated in Table 6. The mPE for our method is smaller than that of Yang et al.’s method [64] on the LOAF test set by 0.097. Similar to the person detection performance, our method realized a noteworthy 0.124 gain in terms of the PEunseen when compared with Yang et al.’s method [64]. Our method outperformed the other methods on PEn, PEm, and PEf because of the accuracy of bounding box detection.
@@ 5 Conclusion
Limitations. We assume that the fisheye images are stereographic projections. However, this assumption may cause a mismatch with the actual distortion. To address this problem, we will extend our method using deep single image camera calibration techniques [4, 54] that estimate the lens distortion. Another promising direction for future work is to use videos for the input. In this paper, we have focused on person detection and localization from an image.